Scalysz:

Given the general identity tan X = sin X/ cos X , which equation relating the acute angles, A and C, of a right ∆ABC is true?

olivia14:

do u have a pic?

Hoodmemes:

Ye need a pic or url of the angles

Scalysz:

A. tan A= sin A/ Sin C B. cos A= tan (90- A)/ sin (90-C) C. sin C= cos A/ tan C D. cos A= tan C E. sin C= cos (90-C)/ tan A

Scalysz:

Scalysz:

sorry theirs the screenshot

olivia14:

its ok, i do edmentum to lol

Hoodmemes:

Ok so tan A = sin A / sin C

Hoodmemes:

and sin C = sin A / tan A = sin A / (sin A / cos A) =

Hoodmemes:

what would it equal?

Scalysz:

I'm pretty sure it's either B or E

Hoodmemes:

but what do u think the answer would be?

Hoodmemes:

btw this is just the second step to finding the answer

Scalysz:

E

Hoodmemes:

why do u think its E?

Hoodmemes:

E is incorrect

Scalysz:

i'm confused

Hoodmemes:

ok so sin C = sin A/tan A = sin A/(sin A / cos A) = cos A sin C =? (btw A/tan is a fraction)

Scalysz:

ok

Hoodmemes:

new answer?

Scalysz:

yea I think I got it now

Hoodmemes:

so what would your final answer be?

Scalysz:

I dont think its B now. Im starting to lean on towards C.

Hoodmemes:

inncorrect

Hoodmemes:

Ok lets break it down again

Scalysz:

so its B?

Hoodmemes:

Your last two options are A and B so give a explination on why u think B or A is correct